# Towards Coordinate- and Dimension-Agnostic Machine Learning for Partial Differential Equations

### v1

@@ Abstract
The machine learning methods for data-driven identification of partial differential equations (PDEs) are typically defined for a given number of spatial dimensions and a choice of coordinates the data have been collected in. This dependence prevents the learned evolution equation from generalizing to other spaces. In this work, we reformulate the problem in terms of coordinate- and dimension-independent representations, paving the way toward what we call “spatially liberated" PDE learning. To this end, we employ a machine learning approach to predict the evolution of scalar field systems expressed in the formalism of exterior calculus, which is coordinate-free and immediately generalizes to arbitrary dimensions by construction. We demonstrate the performance of this approach in the FitzHugh-Nagumo and Barkley reaction-diffusion models, as well as the Patlak-Keller-Segel model informed by in-situ chemotactic bacteria observations. We provide extensive numerical experiments that demonstrate that our approach allows for seamless transitions across various spatial contexts. We show that the field dynamics learned in one space can be used to make accurate predictions in other spaces with different dimensions, coordinate systems, boundary conditions, and curvatures.
@@ 1 Introduction
Machine learning has traditionally been used to identify and model partial differential equations (PDEs) from observational data krischer1993model ; gonzalez1998identification ; blechschmidt2021three ; cuomo2022scientific . In particular, deep learning techniques such as physics-informed machine learning karniadakis2021physics ; raissi2019physics and operator learning lu2021learning ; kovachki2023neural ; fabiani2025randonets , often in combination with symbolic regression techniques brunton2016discovering ; cranmer2023interpretable , can extract spatiotemporal patterns from experimental measurements and (approximately) learn the underlying physical laws. This has been especially useful for describing complex dynamical systems in various scientific and engineering domains, where conventional numerical methods can be computationally expensive kochkov2021machine , or where we only have access to an incomplete description of the system martin2023physics . As a result, machine learning based modeling of PDEs is used on a wide range of scales, from fine-grained descriptions at the particle level lee2023learning to coarse-grained representations in terms of density fields psarellis2024data , providing a data-driven alternative to traditional physics-based derivations.
@@
Real-world data are often expressed in domains and coordinate systems that are determined by the experimental acquisition process. For instance, bacterial chemotaxis can be captured as one-dimensional wave channels adler1966chemotaxis , or visualized as two-dimensional images — top-down with upright phan2020bacterial or bottom-up with inverted microscopy morris2017bacterial . All three experimental settings produce data of the same physical process supported in different geometries and coordinate systems. The conventional machine learning approaches for PDE identification are inherently tied to a predefined spatial framework, preventing them from generalizing the learned physical mechanism to different coordinate systems, geometries, or dimensions. In this work we propose a more flexible approach that would allow learning a dynamical system (a physical mechanistic model) from data in one domain, and then making accurate predictions in another, when the underlying physics is the same. Recent work by Psarellis et al. psarellis2024data has proposed a coordinate-free PDE learning approach using exterior calculus. Their method allows them to learn a PDE independently of the coordinate system, and they show that the learning generalizes to different coordinate system representations of the same domain. We further extend the method to generalize to other domains with different geometries and dimensions.
Our contributions
@@
Inspired by the work by Psarellis et al. psarellis2024data , we propose a machine learning approach that can learn a PDE in one domain and generalize it to different domains and dimensions. Our approach applies to PDEs that describe the evolution of multiple coupled scalar fields. We only consider PDE operators that are invariant with respect to local orthogonal symmetries weyl2015symmetry ; frankel2011geometry ; nakahara2018geometry ; abraham2012manifolds ; villartowards , which is a feature observed in many natural phenomena. For these systems, we show how machine learning can achieve “full spatial liberation,” enabling training in any space and coordinate system, yet being able to make accurate predictions across all others, regardless of coordinates, but also dimensions, boundary conditions, and spatial intrinsic curvatures (i.e. non-vanishing Riemann curvature tensor do1992riemannian ). Fusion of training data obtained at different coordinates/geometries is also enabled.
@@
In our experiments, we use neural networks to identify the FitzHugh-Nagumo bar2003bifurcation and Barkley barkley1991model reaction-diffusion model operators; and a modified Patlak-Keller-Segel model, derived from in-situ observations of chemotactic bacteria phan2020bacterial . We train the models on data generated in a periodic one-dimensional space using Cartesian coordinates, and then demonstrate their ability to generalize to higher-dimensional spaces. These include two- and three-dimensional Euclidean spaces in Cartesian or non-Cartesian coordinates, as well as a two-dimensional submanifold of a sphere in Euclidean space, described using geographic as well as stereographic coordinates. We also consider different boundary conditions in these new spaces, such as a no-flux Neumann boundary.
@@ 2 Coordinate-free Operator Learning Across Dimensions
For most [^1] For a precise statement see page 4 of bryant1999nine . differential equations defined on a Riemannian manifold $\mathcal{M}$ with metric $g$, there exists a coordinate-free mathematical expression that represents them as an exterior differential system (EDS) bryant2013exterior ; mckay2017introduction , in terms of differential forms on $\mathcal{M}$. We describe the language of differential forms and give references in Appendix A. Given an EDS, one can express a particular PDE in coordinate-explicit form after choosing a coordinate system on the manifold $\mathcal{M}$, and transformations of the PDE under diffeomorphisms of $\mathcal{M}$ are governed by the chain rule (pullbacks). Certain EDSs exhibit additional structure that allows us to define them on spaces of varying dimension; we give the following simple example:
Example.
The heat equation of a scalar temperature field can be expressed in a coordinate-free way using differential forms as
$\frac{\partial u}{\partial t}=\star\,d\star du,$ (1)
where $d$ is the exterior derivative and $\star$ is the Hodge star operator.
In addition to being coordinate-free, the heat equation in equation (1) is “domain-agnostic” and “dimension-agnostic” — meaning that the same heat equation can be defined on any manifold $\mathcal{M}$ of any dimension. The dimension and the Riemannian metric appear via the definitions of $d$ and $\star$. In this paper we argue that we can learn the right hand side of an evolution equation in one domain and transfer its learning to other manifolds of different dimensions, as long as the differential equation is defined as a “dimension-agnostic” expression.
@@
Many PDEs of mathematical physics and other fields of science and engineering can be expressed in a coordinate-free manner in terms of their independent variables and their (unknown) fields. Notable examples are the Navier-Stokes wilson2011 equations and, more generally, the equations of continuum mechanics marsden1994 , the Fokker-Planck equation masoliver1987 , the Kuramoto-Sivashinsky equation frankel1987 , as well as many others.
@@
Our work in this paper focuses exclusively on systems of scalar fields (00-forms) governed by PDEs invariant under local orthogonal transformations (rotations and reflections) and under diffeomorphic translations frankel2011geometry ; nakahara2018geometry ; abraham2012manifolds . Such systems arise frequently in nature, as microscopic physical interactions typically exhibit these symmetries; thus, the emergent PDEs inherit them at macroscopic scales. We consider the PDE to involve at most second-order [^2] Higher-order will be considered in future works by extending the invariant features discussed in Eq. (5). spatial-derivatives, in which we adopt the power-counting scheme of effective field theory kaplan2005five ; buchalla2014power with the total differential order of each term defined as the sum of its derivative orders. [^3] This is the power-counting scheme for wave numbers, which is of critical importance when creating an effective field theory for coarse-grained dynamics kaplan2005five ; buchalla2014power . Since each spatial derivative introduces a factor of the wave number (e.g. in Fourier space), higher-order terms are increasingly suppressed at low wave numbers. By explicitly tracking the total derivative order, this method provides a more accurate characterization of the dominant large-scale emergent behavior of the system. We then observe that when we apply these restrictions, the resulting class of PDEs turn out to be domain-free and dimension-free and we can therefore generalize the learning to other domains in other dimensions.
@@
    Given data satisfying an evolution equation, first learn a spatially coordinate-free representation of the evolution operator of the PDE expressed in terms of differential forms,
@@
    Given a new set of initial and/or boundary conditions in a possibly different coordinate system, or even dimension, use a coordinate-free solution of the PDE to evolve the dynamics.
Note that coordinate-free integrators are a classical subject of study in numerical analysis dinesh2000algebraic .
@@
Generators of invariant 0-forms. Given $N$ scalar fields $\quantity{\Psi_{1},\ldots,\Psi_{N}}$ (i.e. 0-forms in a Riemannian manifold $\mathcal{M}$) the set of all possible orthogonally symmetric operators from 00-forms to 00-forms involving up to second-order derivatives can be obtained from:
  • The Laplacians:
    $\star\ d\star d\Psi_{j},$ (2)
    where $j=1,\ldots,N$. There are $N$ of them in total.
  • The inner products:
    $\langle d\Psi_{j},d\Psi_{k}\rangle\equiv\star\left(d\Psi_{j}\wedge\star\ d\Psi_{k}\right),$ (3)
    where $j,k=1,\ldots,N$. Since these are symmetric, there are $N(N+1)/2$ distinct products.
In the above expressions, the Hodge star $\star$ maps $k$-forms to $(n-k)$-forms where $n$ is the dimension of the space. The Laplacians in (2) and the inner products in (3) are with respect to the Riemannian metric of $\mathcal{M}$. These operators can be written explicitly in coordinates using the metric structure of the manifold (see Appendix A).
Coordinate-free learning. Given $\Psi_{1}$,…, $\Psi_{N}$ as above we consider the PDE identification problem
@@
where we assume the $\mathcal{F}_{j}$’s are (possibly non-linear) orthogonally symmetric second order operators defining the right-hand-side of the evolutionary PDE. All second-order, functionally independent scalar invariants (00-forms), can be built from the original scalar fields, their gradient inner products, and their Laplacians, which follows from the first fundamental theorem of classical invariant theory for the orthogonal group weyl1946classical ; spivak1999diffgeoV , which has been extensively used in machine learning villar2021scalars ; gregory2024learning .  [^4] All SO($n$)-scalars can be built locally from derivatives of scalars, the metric (Kronecker delta), and the Levi-Civita symbol via basic arithmetic operations and contractions spivak1999diffgeoV ; for O($n$)-invariance the Levi-Civita symbol is excluded. Our claim comes from applying this theorem to the second-order jet space, in which in this work we restrict to scalar expressions with total order of derivatives at most two. For higher-order, we can list all possible features constructed from the metric, the fields, and its derivatives — up to the desired order — to be used as invariant features.
Remark 1.
Parity invariance under the full orthogonal group O($n$), which extends the rotational symmetries of SO($n$) to include reflections, excludes pseudoscalars. For instance, in dimension $n=2$ it excludes the Jacobian determinant 00-form $\star(d\Psi_{j}\wedge d\Psi_{k})$. Note that this expression does not define a 0-form in dimensions $n\neq 2$.
We construct the invariant features
$\mathcal{B}=([\Psi_{j}]_{j=1}^{N},\,[\star\ d\star d\Psi_{j}]_{j=1}^{N},\,[\langle d\Psi_{j},d\Psi_{k}\rangle]_{1\leq j\leq k\leq N})=:\left(b_{1},\ldots,b_{\frac{N(N+5)}{2}}\right)$ (5)
@@
to use them as input features for neural networks ^$\hat{\mathcal{F}}_{j}(\mathcal{B})$ ($j=1,\ldots,N$) that learn the right-hand-sides of (4). Namely, it learns functions ^$\hat{\mathcal{F}}_{j}(\mathcal{B})$ such that the integration of the corresponding PDE via a coordinate-free integrator approximates the training data. A coordinate-free integrator in this context is one that performs the integration of the PDE in terms of the features $b_{i}$. Note that this approach can be seen as a form of canonicalization (see kaba2023equivariance ; olver2003moving ), where the input data is transformed into a set of canonical invariants (here $\mathcal{B}$) to perform the learning.
@@
Any-dimensional coordinate-free learning. The Laplacians and inner products are defined as functions of the input scalar fields for any domain –any Riemannian manifold $(\mathcal{M},g)$ of any dimension $n$. Therefore $\mathcal{B}$ is defined via a domain- and dimension-independent description. The ^$\hat{\mathcal{F}}_{j}$’s are defined as functions of $N(N+5)/2$ features, $(b_{1},\ldots,b_{N(N+5)/2})$, independently of the size of the domain of the $b_{i}$’s (and thus independently of the dimension of the domain of the input 0-forms $\Psi_{j}$’s). Therefore the ^$\hat{\mathcal{F}}_{j}$ can be used to predict the dynamics of a family of PDEs defined in any input space.
This principle, of using a fixed set of parameters to define models in spaces of any size, is common in machine learning. In particular, these ideas appear in equivariant machine learning, where models such as graph neural networks use a fixed set of parameters to define functions on input graphs of any size ruiz2020graphon ; maskey2023transferability ; velasco2024graph . The idea in the context of this work is illustrated in Figure 1. The theoretical treatment of the topic in the context of machine learning is pioneered in recent work by Levin and Díaz levin2024any , using tools from the field of representation stability farb2014representation .
3 Examples
In this work, we apply the proposed methodology to three PDEs that can be expressed in a coordinate-free and dimension-free form, each involving two coupled scalar fields (thus $N=2$ in Eq. (4) and Eq. (5)).
3.1 The FitzHugh-Nagumo System

### v2

@@ Abstract
Machine learning methods for data-driven identification of partial differential equations (PDEs) are typically defined for a fixed number of spatial dimensions and a particular choice of coordinates in which the data have been collected. This dependence prevents the learned equation from generalizing to other spaces. In this work, we reformulate the problem in terms of coordinate- and dimension-independent representations, paving the way toward what we might call “spatially liberated" PDE learning. In this work, we propose an approach to learning differential equations by expressing them in a way that is independent of the coordinate system and even the underlying manifold where the equation is defined. This allows us to learn a differential equation in low-dimensional spaces and generalize to higher-dimensional spaces with different geometric properties. We provide extensive numerical experiments that demonstrate that our approach allows for seamless transitions across various geometric contexts. We show that the dynamics learned in one space can be used to make accurate predictions in other spaces with different dimensions, coordinate systems, boundary conditions, and curvatures.
Preprint — October 2025
@@ 1 Introduction
Machine learning has traditionally been used to identify and model partial differential equations (PDEs) from observational data (Krischer et al., 1993; González-García et al., 1998; Blechschmidt & Ernst, 2021; Cuomo et al., 2022). In particular, deep learning techniques such as physics-informed machine learning (Karniadakis et al., 2021; Raissi et al., 2019) and operator learning (Lu et al., 2021; Kovachki et al., 2023; Fabiani et al., 2025), often in combination with symbolic regression techniques (Brunton et al., 2016; Cranmer, 2023), can extract spatiotemporal patterns from experimental measurements and (approximately) infer the underlying physical laws. This has been especially useful for describing complex dynamical systems in various scientific and engineering domains, where conventional numerical methods can be computationally expensive (Kochkov et al., 2021), or where we only have access to an incomplete description of the system (Martin-Linares et al., 2023). As a result, learning-based PDE models are used on a wide range of scales, from fine-grained descriptions at the particle level (Lee et al., 2023) to coarse-grained representations in terms of density fields (Psarellis et al., 2024), providing a data-driven alternative to traditional physics-based derivations.
@@
Real-world data are often expressed in domains and coordinate systems that are determined by the experimental acquisition process. For instance, bacterial chemotaxis can be captured as waves in effectively one-dimensional channels (Adler, 1966), or visualized as two-dimensional images — top-down, with upright (Phan et al., 2020) or bottom-up, with inverted microscopy (Morris et al., 2017). All three experimental settings produce data of the same physical process supported in different geometries and coordinate systems. The conventional machine learning approaches for PDE identification are inherently tied to a predefined spatial framework, preventing them from generalizing the learned physical mechanism to different coordinate systems, geometries, or dimensions. The recent work of Psarellis et al. (2024) has proposed a coordinate-independent PDE learning approach using exterior calculus. Their method allows them to learn a time evolution scalar PDE independently of the coordinate system, and they show that the learning generalizes to different coordinate system representations of the same domain.
In this work we push these ideas substantially further, and show that we can learn PDEs not just in a coordinate-independent way but also independently of the data domain, dimension, and geometry.
We propose a framework that allows us to learn a dynamical system (a physical mechanistic model) from data in one domain, and then make accurate predictions in another, when the underlying physics is the same.
We are able to do this by combining ideas from the coordinate-free learning of time evolution scalar PDEs from Psarellis et al. (2024), and by recent work in any-dimensional learning that proposes a theoretical framework for machine learning models that are defined on a fixed set of parameters but can be trained and evaluated in spaces of arbitrary size (Levin & Díaz, 2024; Levin et al., 2025).
The most widely used any-dimensional models are arguably graph neural networks (GNNs), where a model defined on a fixed set of parameters can be trained and evaluated on graphs of any size (and GNNs are completely agnostic to the graph sizes). In order to have a sufficiently expressive machine learning model with a fixed number of parameters while the dimension of the input space grows arbitrarily, one typically relies on some form of symmetry constraint (this is mathematically formalized via representation stability (Farb, 2014)). For example, graph neural networks are equivariant with respect to node permutations.
@@
In this work we learn PDE operators that are invariant with respect to local orthogonal symmetries (Weyl, 2015; Frankel, 2011; Nakahara, 2018; Abraham et al., 2012; Villar et al., 2023), which is a feature observed in many natural phenomena. For these systems, we show how machine learning can achieve “full spatial liberation”, enabling training in any space and coordinates, yet being able to make accurate predictions across all others, regardless of coordinates, but also dimensions, boundary conditions, and spatial intrinsic curvatures (i.e. Riemann curvature tensors (Do Carmo & Flaherty Francis, 1992)). Fusion of training data obtained at different coordinates/geometries is also enabled.
@@
In our experiments, we use neural networks to identify time evolution PDEs of coupled scalar fields, general scalar PDEs, and vector PDEs in a coordinate-free and dimension-free way. In particular, we apply our models to the FitzHugh-Nagumo (Bär et al., 2003) and Barkley (Barkley, 1991) reaction-diffusion model operators; and a modified Patlak-Keller-Segel model, derived from in-situ observations of chemotactic bacteria (Phan et al., 2020). We also illustrate our method on the Helmholtz equation, and the incompressible Navier-Stokes equation. Most of our models are trained on data generated in a periodic one-dimensional space using Cartesian coordinates, and we then demonstrate their ability to generalize to higher-dimensional spaces. These include two- and three-dimensional Euclidean spaces in Cartesian or non-Cartesian coordinates, as well as a two-dimensional submanifold of a sphere in Euclidean space, described using geographic as well as stereographic coordinates. We also consider different boundary conditions in these new spaces, such as a no-flux Neumann boundary.
@@ 2 Coordinate-free Operator Learning Across Dimensions
Many differential equations can be formulated in a coordinate-free and domain-free mathematical expression. For example, the heat equation written as $\frac{\partial u}{\partial t}=\sum_{i=1}^{n}\frac{\delta^{2}u}{\delta x_{i}^{2}}$ is not coordinate free, since it is expressed in terms of local coordinates. However the heat equation expressed as $\frac{\partial u}{\partial t}=\Delta u$ (where $\Delta$ is the Laplacian operator of the Riemannian manifold $\mathcal{M}$ where $u$ is defined) is not only coordinate-free, since the Laplacian $\Delta$ is an intrinsic operator of $\mathcal{M}$, but also it is domain-free, since the Laplacian exists for every Riemannian manifold. Therefore, the equation is well defined and can be evaluated for manifolds of any dimension. This procedure can be formalized in terms of differential forms and exterior differential systems (see Appendix A). In this paper we argue that we can learn the PDE in one domain and transfer its learning to other manifolds of different dimensions, as long as the differential equation is defined as a “dimension-agnostic” expression.
@@
PDEs coming from mathematical physics and other fields of science and engineering often can be expressed in a coordinate-free manner in terms of their independent variables and their (unknown) fields. Notable examples are the Navier-Stokes (Wilson, 2011) and, more generally, the equations of continuum mechanics (Marsden & Hughes, 1994), the Fokker-Planck equation (Masoliver et al., 1987), the Kuramoto-Sivashinsky equation (Frankel & Sivashinsky, 1987), and many others.
@@
In this paper we consider systems of scalar and vector PDEs. All of them governed by equations that are invariant under local orthogonal transformations (rotations and reflections) and under diffeomorphic translations (Frankel, 2011; Nakahara, 2018; Abraham et al., 2012). Such systems arise frequently in nature, as microscopic physical interactions typically exhibit these symmetries; thus, the emergent PDEs inherit them at macroscopic scales. We consider the PDE to involve at most second-order [^1] Higher-order will be considered in future works by extending the invariant features discussed in Eq. (1). spatial-derivatives, in which we adopt the power-counting scheme of effective field theory (Kaplan, 2005; Buchalla et al., 2014) with the total differential order of each term defined as the sum of its derivative orders. [^2] This is the power-counting scheme for wave numbers, which is of critical importance when creating an effective field theory for coarse-grained dynamics. Since each spatial derivative introduces a factor of the wave number (e.g. in Fourier space), higher-order terms are increasingly suppressed at low wave numbers. By explicitly tracking the total derivative order, this method provides a more accurate characterization of the dominant large-scale emergent behavior of the system. We then observe that when we apply these restrictions, the resulting class of PDEs turn out to be domain-free and dimension-free and we can therefore generalize the learning to other domains in other dimensions.
@@
    Given data satisfying a partial differential equation, first learn a spatially coordinate-free representation of the PDE.
@@
    Given a new set of initial and/or boundary conditions in a possibly different coordinate system, or even dimension, use a coordinate-free solver of the PDE to evolve the dynamics.
Coordinate-free integrators are a classical subject of study in numerical analysis (Dinesh et al., 2000).
@@
Generators of invariant fields. Given $N$ scalar fields $\quantity{\Psi_{1},\ldots,\Psi_{N}}$ (i.e. 0-forms on a Riemannian manifold $\mathcal{M}$) the set of all possible orthogonally symmetric operators from 0-forms to 0-forms involving up to second-order derivatives can be obtained from [^3] This follows from the first fundamental theorem of classical invariant theory for the orthogonal group (Weyl, 1946; Spivak, 1999), which has been extensively used in machine learning (Villar et al., 2021; Gregory et al., 2024). All SO($n$)-scalars can be built locally from derivatives of scalars, the metric (Kronecker delta), and the Levi-Civita symbol via basic arithmetic operations and contractions (Spivak, 1999); for O($n$)-invariance the Levi-Civita symbol is excluded. Our claim comes from applying this theorem to the second-order jet space, in which in this work we restrict to scalar expressions with total order of derivatives at most two. For higher-order, we can list all possible features constructed from the metric, the fields, and its derivatives — up to the desired order — to be used as invariant features. :
  • The fields $\Psi_{j}$ where $j=1,\ldots,N$.
  • The Laplacians $\Delta\Psi_{j}$, where $j=1,\ldots,N$.
  • The inner products of the gradients $\langle\gradient\Psi_{j},\gradient\Psi_{k}\rangle$, where $j,k=1,\ldots,N$. Since these are symmetric, there are $N(N+1)/2$ distinct products.
The Laplacians and the inner products of the gradients above are defined with respect to the Riemannian metric of $\mathcal{M}$, and can be expressed in coordinate-free form using the metric structure of the manifold (see Appendix A).
This yields a basis of invariant features $\mathcal{B}(\Psi_{1},\ldots,\Psi_{N})$ (or $\mathcal{B}$ when the input is clear by context),
$\mathcal{B}(\Psi_{1},\ldots,\Psi_{N})=([\Psi_{j}]_{j=1}^{N},\,[\Delta\Psi_{j}]_{j=1}^{N},\,[\langle\nabla\Psi_{j},\nabla\Psi_{k}\rangle]_{1\leq j\leq k\leq N})=:\left(b_{1},\ldots,b_{\frac{N(N+5)}{2}}\right).$ (1)
Coordinate-free learning of evolution PDEs. Given $\Psi_{1}$,…, $\Psi_{N}$ as above we consider the PDE identification problem
@@
where we assume the $\mathcal{F}_{j}$’s are (possibly non-linear) orthogonally symmetric second order operators defining the right-hand-side of the evolutionary PDE. We consider neural networks that learn $\hat{\mathcal{F}}_{j}(\mathcal{B})$ as functions of the features (1), such that the integration of the corresponding PDE via a coordinate-free integrator approximates the training data. A coordinate-free integrator in this context is one that performs the integration of the PDE in terms of the features $b_{i}$. Note that this approach can be seen as a form of canonicalization (see Kaba et al. (2023); Olver (2003)), where the input data is transformed into a set of canonical invariants (here $\mathcal{B}$) to perform the learning.
Coordinate-free learning of general scalar PDEs. The time evolution PDEs we consider here can be expressed as $\frac{\partial u}{\partial t}=\mathcal{F}(\mathcal{B}(u))$ where $\frac{\partial u}{\partial t}$ is the independent variable. A general PDE is an implicit algebraic constraint. We therefore do not a priori know which are the independent or dependent variables. Instead, we can utilize manifold learning methods, linear (e.g. PCA) for linear PDEs, nonlinear (e.g. autoencoders or kernel-PCA) for nonlinear PDEs, to discover the implicit constraint within the coordinate-free basis. In this way, the PDE operator is represented by a latent model constraining the relationship between the variables in $\mathcal{B}$. After this is learned, it can be ‘extracted’ symbolically using symbolic regression, or used implicitly by numerical solvers. Of course, there is the caveat that, when the PDE is nonlinear, the constraint learned is only guaranteed to be accurate in the region adequately sampled by the training data.
The workflow for a general PDE is as follows:
  1. 1.
    Project the data satisfying the PDE, $u$, to the coordinate- and dimension-free basis $\mathcal{B}(u)$.
  2. 2.
    Use manifold-learning to parametrize the submanifold defined by the data in terms of $\mathcal{B}(u)$.
  3. 3.
    Use the trained model as a representation of the operator on different domains, or extract a symbolic representation of the operator using symbolic regression, before transferring to another domain or dimension.
We implement this method for the Helmholtz equation in Section 3.2. We remark that, in order to solve a higher-dimensional problem, we must have information about the ‘well-posedness’ of that problem. For example, the harmonic oscillator needs only two scalar initial/boundary conditions to be solved, whereas the Helmholtz equation needs continua of boundary conditions to be well-posed. This information is not encoded in our coordinate-free representation of the operator in general. However, for time evolution problems, the required initial/boundary conditions are the same across other domains and dimensions.
Extension to vector PDEs. Our proposed approach can be extended to vector PDEs, or more generally coupled scalar and vector field PDEs, by computing an extended coordinate-free feature basis $\tilde{\mathcal{B}}$, which includes scalar features $\mathcal{S}$ and vector features $\mathcal{V}$. Given this basis, a general class of coupled scalar-vector PDEs takes the form:
$\mathcal{F}(\tilde{\mathcal{B}})=\sum_{i=1}^{|\mathcal{V}|}f_{i}\quantity({\quantity{s_{j}}_{j=1}^{|\mathcal{S}|}})v_{i}=0$ (3)
where $f_{i}:\mathcal{S}\to\mathbb{R}$ are arbitrary nonlinear functions. This equation matches the parameterization of equivariant functions to actions by the orthogonal group from Villar et al. (2021). We bound the degree of the differentials in the basis by $2$ in our work; this can, however, be relaxed based on the problem at hand.
The methodology for constructing coordinate-free features from a scalar field $s$ remains the same as in (1). In particular, the scalar features up to order-$2$ will be $\quantity{s,\Delta s}\in\mathcal{S}$, and the sole vector feature will be $\gradient{s}$. For a vector feature $v$ we will have $\expectationvalue{v,v}\in\mathcal{S}$ and $v,\Delta v\in\mathcal{V}$, where the latter operator is the component-wise Laplacian. Finally, we include terms of the form $\expectationvalue{v,\gradient\cdot v}\in\mathcal{V}$, which can be formulated in a cordinate-free manner in terms of the Lie derivative, and terms of the form $\divergence{v}\in\mathcal{S}$, denoting divergence. The remaining elements of the library are then computed using the inner product, divergence, and Lie-derivative operations, all of which have coordinate-free formulations in terms of exterior algebra.
@@
Any-dimensional coordinate-free learning. The Laplacians and inner products are defined as functions of the input scalar fields for any domain –any Riemannian manifold $(\mathcal{M},g)$ of any dimension $n$. Therefore $\mathcal{B}$ (or analogously $\tilde{B}$ in the vector PDE case) is defined via a domain- and dimension-independent description. The learned PDEs are defined as functions of $N(N+5)/2$ features, $(b_{1},\ldots,b_{N(N+5)/2})$, independently of the size of the domain of the $b_{i}$’s (and thus independently of the dimension of the domain of the inputs).
3 Computational Experiments
3.1 Time-Evolution Systems
In this section, we apply the proposed methodology to three time-evolution PDEs: the FitzHugh-Nagumo, the Barkley, and a modified Patlak-Keller-Segel Systems defined below. Each can be expressed in a coordinate-free and dimension-free form (Appendix C), involving two coupled scalar fields (thus $N=2$ in Eq. (2) and Eq. (1)). We present selected computational experiments in the main text, while extensive additional studies are included in the appendices. See Appendices EL for the experimental setup details and quantitative evaluation of the results.